Reduce Boolean equation X=BC+ABC+BC+ABC to a minimum SOP form. A/
Answer #1

Ans)

Given

X=BC+A'B'C+BC'+AB'C

X=BC+BC'+(A'+A)B'C using rule that Z'+Z=1

X=BC+BC'+B'C

now using Y+Y=Y

X=BC+BC'+BC+B'C

X=B(C+C')+(B+B')C

\mathbf{X=B+C} is the answer

========================